I am stuck on this problem!
At one p.m., 1/5 of the geese fly away. At two p.m., 1/8 of the geese fly away. At three p.m., three times as many geeseflew away as one p.m. There are now 28 geese left at the pond. How many geese were in the original flock?
I tried this equation: X-1/5x-1/8x-3/5x = 28, but get 373.333 or 373 and 1/3rd geese? What am i doing wrong? Thanks

Found 2 solutions by stanbon, ptaylor:
Answer by stanbon(75887)

At one p.m., 1/5 of the geese fly away. At two p.m., 1/8 of the geese fly away. At three p.m., three times as many geeseflew away as one p.m. There are now 28 geese left at the pond. How many geese were in the original flock?
----------------
Let the number of geese in the original flock be "x".
EQUATION:
x - (1/5)x = (4/5)x
That is the number left after 1 PM.
------------------------
So at 2 PM (1/8)(4/5)x fly away
That leaves (7/8)(4/5)x = (7/10) x after 2 PM
-----------------------------
At 3 PM you lose (3/5)x geese
leaving you with (7/10)x - (3/5)x = (1/10)x
-------------------------
EQUATION:
(1/10)x = 28
x = 280 geese (# you started with)
================
Cheers,
Stan H.

Answer by ptaylor(2198)

Let x=number of geese in the original flock
At one pm, x/5 fly away
and at one pm, we have left x-x/5=4x/5 geese
At two pm (1/8)(4x/5) or(4x/40) or (x/10) fly away-------------Now I'm assuming that you mean 1/8 of the geese that are left
At two pm, (4x/5)-(x/10) or (8x/10)-(x/10) or 7x/10 geese are left
At three pm, 3*(x/5) flew away that means that
At three pm, (7x/10)-(3x/5) or (7x/10)-6x/10) or x/10 geese are left and we are told that this equals 28.--------------So:
x/10=28 multiply both sides by 10
x=280-----------------------------------number of geese in the original flock
CK
at one pm, 280/5 or 56 geese flew away leaving 280-56 or 224
at two pm 224/8 or 28 geese flew away leaving 224-28 or 196
At three pm 3*56 or 168 geese flew away leaving 196-168 or 28
Hope this helps---ptaylor